# Decoding Financial Health in Kenya’s Medical Insurance Sector: A Data-Driven Cluster Analysis

### v1

@@ Abstract
This study examines insurance companies’ financial performance and reporting trends within the medical sector using advanced clustering techniques to identify distinct patterns. Four clusters were identified by analyzing financial ratios and time series data, each representing unique financial performance and reporting consistency combinations. Dynamic Time Warping (DTW) and KMeans clustering were employed to capture temporal variations and uncover key insights into company behaviors. The findings reveal that resilient performers consistently report and have financial stability, making them reliable options for policyholders. In contrast, clusters of underperforming companies and those with reporting gaps highlight operational challenges and issues related to data consistency. These insights emphasize the importance of transparency and timely reporting to ensure the sector’s resilience. This study contributes to the literature by integrating time series analysis into financial clustering, offering practical recommendations for improving data governance and financial stability in the insurance sector. Future research could further investigate non-financial indicators and explore alternative clustering methods to provide a deeper understanding of performance dynamics.
@@ 1   Introduction
Sustainability in the insurance sector is closely related to financial performance. Effective risk management enhances operational efficiency, enabling insurers to meet their obligations and maintain policyholder confidence. Given the sector’s critical role in safeguarding financial and health security, there is a pressing need to evaluate the performance of insurance companies and identify factors influencing their competitiveness. This study addresses this gap by clustering insurance firms based on financial ratios derived from open data. Temporal clustering of quarterly time series data will facilitate comparisons over time, shedding light on the relative performance of insurers.
The findings of this study have implications for multiple stakeholders. Policymakers can leverage the insights to improve sectoral efficiency through regulatory reforms, insurers can identify strategic opportunities for improvement, and policyholders can make informed decisions when selecting providers. The study also contributes to the broader discourse on the intersection of financial performance, risk management, and sustainability in the insurance sector.
This paper is structured as follows: Section 2 provides an overview of the relevant literature. Section 3 details the methodology, including the research design, data sources, financial ratio selection, application of a temporal feature fusion, clustering analysis, and cluster evaluation metrics. Section 4 reports the findings, followed by Section 5, which discusses the key results, addresses the study’s limitations, and proposes avenues for future research. Finally, Section 6 concludes the paper. All analyses were performed in Python, using a variety of statistical and machine-learning libraries.
@@ 2   Literature Review
Health insurance, particularly in developing countries like Kenya, is a critical element of the healthcare system, providing financial protection against the ever-increasing costs of medical care. The private sector, which plays a dominant role in Kenya’s health insurance market, faces substantial challenges in extending coverage to low-income populations, especially in rural areas. Despite the growing importance of health insurance in Kenya, the penetration rate remains relatively low. According to the Kenya Insurance Regulatory Authority (IRA), the penetration rate of health insurance in Kenya was recorded at only 2.25% in 2021, with significant disparities between urban and rural regions [1]. The insurance density was reported at 4777, highlighting the relatively low participation of the population in private health insurance programs. These challenges are compounded by the reliance on out-of-pocket payments, further straining the financial security of individuals and families.
Much of the existing literature focuses on insurance providers’ financial health, operational efficiency, and market dynamics, primarily in developed markets. However, there is a noticeable gap in studies specifically addressing health insurers’ financial performance and operational efficiency in sub-Saharan Africa, particularly in Kenya. The few studies conducted in Kenya, such as [5, 17, 19, 20, 21], have focused mainly on individual aspects of financial performance but have not provided a holistic analysis of the interrelationships between key financial ratios or the general health of the insurance market.
A paper by the authors in [27] is one of the few studies exploring financial success drivers within the Kenyan insurance industry. This study investigates the factors influencing the economic performance of insurance firms using a sample of 37 general insurers and 16 life insurers from 2009 to 2018. The author employs panel data methods to analyze determinants such as insurers’ size, leverage, and age, including pooled OLS, fixed effects, and random effects models. The results show a positive relationship between insurer size and financial performance, while higher leverage was positively correlated with better performance. This study provides valuable information on the broader drivers of economic performance but does not explicitly explore how different types of insurers, such as health insurers, can behave differently concerning financial ratios.
Another study [28] examines the effects of various insurance risks, such as credit risk, liquidity risk, solvency risk, reinsurance risk, and underwriting risk, on the financial performance of insurance companies in Kenya. This study uses a census approach of the 53 licensed insurance companies in Kenya and uses panel data analysis techniques to explore the relationship between risk factors and financial performance. The study finds that credit risk, liquidity risk, solvency risk, and underwriting risk all negatively impact financial performance, while reinsurance risk had a positive insignificant effect.
At an international level, several studies have used financial ratios to assess companies’ profitability, efficiency, solvency, and overall market competitiveness in various industries. A survey by [42] discusses the use of financial ratios as sequences that can be clustered into groups for better evaluation. Using fuzzy clustering methods, the study explores how financial ratios can be grouped based on their similarities, helping to identify common characteristics among companies with similar financial performance.
Similarly, [9] used cluster analysis to classify companies based on their financial strategies. This study analyzed financial coefficients representing strategies for the 2030 largest non-financial companies from 2006 to 2018. Seven clusters were identified, highlighting differences in strategy between high-tech and basic industries. In particular, high-tech companies exhibited distinct strategic patterns compared to those in essential industries. The analysis also revealed a growing prevalence of risky strategies over time. The meaningful interpretation of the clustering results validated the robustness and consistency of the findings.
Herman et al. [15] further explore cluster analysis in the context of Hungarian and Romanian food retail companies, using financial ratios such as Return on Sales (ROS), Return on Assets (ROA), and Return on Equity (ROE). This study emphasizes the need for careful selection of clustering methods, as different approaches can produce different results. The findings indicate that the chosen method for grouping—whether K-Means or K-Medoid—can significantly influence the classification of companies. However, this study is limited by its focus on retail sectors, which differs from the focus on health insurance.
Qian [35] introduced a financial K-means algorithm for classifying listed companies based on financial ratios. This study, which applies cluster analysis to companies in Zhejiang province, highlights the potential of financial ratios in determining loan eligibility. Although the methodology is relevant to financial classification, it does not explore the dynamic behavior of companies over time, nor does it focus on the health insurance sector.
Despite these contributions, there is a noticeable gap in the literature concerning the use of time-series financial ratios for clustering insurance companies, especially in the context of developing countries. Many studies, such as those by [15, 35, 42], rely on static, one-point-in-time data, making it challenging to assess how financial performance evolves. Moreover, there is limited research on using more advanced methodologies, such as dimension reduction and dynamic time-warping clustering algorithms, which could provide deeper insights into the temporal dynamics of financial ratios. Another critical gap in the existing literature is the lack of research on applying financial ratios in the Kenyan insurance market. Studies such as those by [19, 20, 21] focus on the claim incurred ratio for some selected general insurance companies but do not provide a comprehensive view of how this and other ratios interlink and contribute to the overall financial health of insurance providers. Furthermore, the role of the Claims Payout Ratio, which is crucial for understanding an insurer’s efficiency in handling claims, has not been adequately addressed in Kenya’s health insurance market.
@@ 3.1   Research Design
The research design integrates statistical and machine learning methods to extract meaningful insights from the data. Techniques such as Long Short-Term Memory (LSTM) for temporal feature fusion, Dynamic Time Warping (DTW) for analyzing temporal trends, and K-means clustering for identifying patterns were employed to examine the financial performance of insurance providers. These methods were implemented using Python’s object-oriented programming framework, taking advantage of Scikit Learn and clustering algorithms. In addition, several open-source libraries were used to visualize financial ratios, trends, and other relevant metrics. The complete code for this study is publicly accessible on GitHub, promoting further development of the analysis framework.
@@ 3.2   Data
These data, covering multiple fiscal years, were sourced directly from the IRA website [1], ensuring both accuracy and adherence to regulatory standards.
No additional data collection methods, such as surveys or interviews, were utilized. To maintain data integrity, any missing values for companies were systematically recorded as zero. Furthermore, negative values for Gross Premium Income, Claims Paid, and Claims Incurred were adjusted to zero, as such figures are unrealistic within the insurance industry. Comprehensive data preprocessing was conducted to ensure consistency and quality, making the dataset well-suited for robust statistical analysis.
@@ 3.3   Financial Ratios
This study employs key financial ratios to assess the performance and financial health of health insurance providers in Kenya. These ratios are derived from the data metrics presented in Table 1 and provide insights into operational efficiency, profitability, and market position.
A critical measure in this analysis is the Expense Ratio, which reflects the proportion of an insurer’s operational costs relative to its income. It is calculated as the ratio of Expenses to Gross Premium Income (GPI). The Expenses are estimated using the formula:
@@
In addition to the Expense Ratio, the study evaluates other financial ratios, including the Market Share, Claims Paid Ratio, Claims Incurred Ratio, Claims Payout Ratio, Underwriting Profit Ratio, and the Combined Ratio, as detailed in Table 2.
To address potential distortions caused by extreme values, predefined thresholds were applied to cap certain ratios, ensuring a more accurate and reliable analysis. Specifically:
  • Claims Paid Ratio, Claims Incurred Ratio, and Claims Payout Ratio were capped at 500% to mitigate the impact of exceptionally high values from a few companies.
  • Underwriting Profit Ratio was capped at 300% to control for unusually high profit margins.
  • Expense Ratio and Combined Ratio were capped at 800% to exclude extreme operational and loss figures.
These caps, outlined in Table 2, were implemented to minimize the influence of outliers, which may result from reporting errors or exceptional cases.
The financial ratios were normalized using the StandardScaler, standardizing each feature to have a mean of 0 and a standard deviation of 1. A within-company scaling approach was used, where scaling parameters were computed separately for each company to preserve unique statistical characteristics.
Ratio             Formula Global Range Threshold
Market Share PremiumsTotal Market $\frac{\text{Premiums}}{\text{Total Market Premiums}}\times 100$            5% - 30%            N/A
Claims Paid Ratio Claims PaidGross Premium $\frac{\text{Claims Paid}}{\text{Gross Premium Income}}\times 100$            50% - 80%            500%
Loss Ratio Claims IncurredGross Premium $\frac{\text{Claims Incurred}}{\text{Gross Premium Income}}\times 100$            60% - 75%            500%
Underwriting Profit Ratio Underwriting ProfitGross Premium $\frac{\text{Underwriting Profit}}{\text{Gross Premium Income}}\times 100$            5% - 15%            300%
Expense Ratio ExpensesGross Premium $\frac{\text{Expenses}}{\text{Gross Premium Income}}\times 100$            10% - 25%            800%
@@
Claims Payout Ratio Claims PaidClaims $\frac{\text{Claims Paid}}{\text{Claims Incurred}}\times 100$            60% - 90%            500%
Table 2: Key Financial Ratios Used to Evaluate Insurance Providers. Threshold is introduced to limit extreme values and mitigate the impact of outliers. The global range represents typical industry standards for each ratio, indicating the desired financial performance for insurance companies.
@@ 3.4   Long Short-Term Memory (LSTM)
The Long Short-Term Memory (LSTM) model is a specialized recurrent neural network architecture designed to effectively capture temporal dependencies in sequential data. It achieves this through memory blocks, which retain relevant information over time while selectively discarding irrelevant data using gating mechanisms. These mechanisms include input, forget, and output gates, enabling the model to learn both short-term and long-term dependencies [12, 41, 25, 37].
In this study, the Long Short-Term Memory (LSTM) model was utilized to perform temporal feature fusion and reduce the dimensionality of financial data. The input data tensor, denoted as $\mathbf{R}\in\mathbb{R}^{N\times J\times F}$, represents the financial records of $N=28$ insurance companies over $J=41$ time points, with $F=7$ financial ratios recorded at each time step. The objective of the LSTM model is to reduce the feature dimension from $F=7$ to a single temporal dimension ($F^{\prime}=1$), resulting in a reduced data tensor $\mathbf{Z}\in\mathbb{R}^{N\times J\times 1}$, while retaining the temporal dependencies within the data. Figure 2 displays the architecture of a vanilla LSTM block [41]. Each LSTM memory block operates as follows:
@@
The LSTM processes each time-series data vector $\mathbf{r}_{i,j}=[r_{i,j,1},r_{i,j,2},\dots,r_{i,j,F}]^{T}\in\mathbb{R}^{F}$, which corresponds to the $F$-dimensional financial ratios of company $i$ at time $j$. The model outputs a reduced representation $\mathbf{z}_{i,j}\in\mathbb{R}^{1}$, effectively mapping the input tensor $\mathbf{R}$ to the output tensor $\mathbf{Z}$. The LSTM memory block processes the input $\mathbf{r}_{i,j}\in\mathbb{R}^{F}$ alongside the previous hidden state $\mathbf{y}_{i,j-1}\in\mathbb{R}^{y}$ and the cell state $\mathbf{c}_{i,j-1}\in\mathbb{R}^{y}$, where $y$ denotes the number of hidden units in the LSTM.
The block input $\mathbf{g}_{i,j}\in\mathbb{R}^{y}$ combines the current input $\mathbf{r}_{i,j}$ and the previous hidden state $\mathbf{y}_{i,j-1}$ as:
where $\mathbf{W}_{g}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{g}\in\mathbb{R}^{y\times y}$, and $\mathbf{b}_{g}\in\mathbb{R}^{y}$ are learnable parameters. The input gate $\mathbf{i}_{i,j}\in\mathbb{R}^{y}$ determines the extent to which $\mathbf{g}_{i,j}$ influences the cell state:
where $\sigma$ is the sigmoid activation function, and $\mathbf{W}_{i}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{i}\in\mathbb{R}^{y\times y}$, $\mathbf{b}_{i}\in\mathbb{R}^{y}$ are learnable parameters, and $\mathbf{p}_{i}\in\mathbb{R}^{y}$ represents the peephole weights associated with the input gate. The forget gate $\mathbf{f}_{i,j}\in\mathbb{R}^{y}$ determines which information from the previous cell state is retained:
where $\mathbf{W}_{f}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{f}\in\mathbb{R}^{y\times y}$, $\mathbf{b}_{f}\in\mathbb{R}^{y}$, and $\mathbf{p}_{f}\in\mathbb{R}^{y}$ are learnable parameters. The cell state $\mathbf{c}_{i,j}\in\mathbb{R}^{y}$ is updated as:
$\mathbf{c}_{i,j}=\mathbf{f}_{i,j}\odot\mathbf{c}_{i,j-1}+\mathbf{i}_{i,j}\odot\mathbf{g}_{i,j},$
where $\odot$ denotes element-wise multiplication. The output gate $\mathbf{o}_{i,j}\in\mathbb{R}^{y}$ controls the information passed to the final output:
where $\mathbf{W}_{o}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{o}\in\mathbb{R}^{y\times y}$, $\mathbf{b}_{o}\in\mathbb{R}^{y}$, and $\mathbf{p}_{o}\in\mathbb{R}^{y}$ are the peephole weights and learnable parameters. The updated hidden state $\mathbf{y}_{i,j}\in\mathbb{R}^{y}$ is computed as:
$\mathbf{y}_{i,j}=\tanh(\mathbf{c}_{i,j})\odot\mathbf{o}_{i,j}.$
The LSTM output $\mathbf{z}_{i,j}\in\mathbb{R}^{1}$ is obtained by applying a fully connected dense layer to the hidden state:
$\mathbf{z}_{i,j}=\mathbf{W}_{z}\mathbf{y}_{i,j}+\mathbf{b}_{z},$
where $\mathbf{W}_{z}\in\mathbb{R}^{1\times y}$ and $\mathbf{b}_{z}\in\mathbb{R}^{1}$ are learnable parameters. The overall transformation performed by the LSTM network can be expressed as:
$\mathbf{Z}=\psi_{\text{LSTM}}(\mathbf{R}),$
where $\psi_{\text{LSTM}}$ denotes the LSTM’s mapping from the input space $\mathbb{R}^{N\times J\times F}$ to the output space $\mathbb{R}^{N\times J\times 1}$.
@@
where $\hat{\mathbf{R}}$ is the reconstructed input tensor. This reduced representation $\mathbf{Z}$ was subsequently used for clustering analysis, as detailed in Section 3.5.
Figure 3: LSTM Heatmap. The heatmap displays the reduced time series for each company after applying temporal feature fusion using LSTM, which transforms the data from $\mathbb{R}^{7}$ to $\mathbb{R}^{1}$. Color intensity indicates the magnitude of the reduced data values, with darker shades representing higher values. The x-axis represents time points, and the y-axis lists the companies, which will be clustered based on this dataset.

### v2

@@ Abstract
This study explores the financial performance and reporting trends of insurance companies in the medical sector using advanced clustering techniques to identify distinct patterns. By analyzing financial ratios and time series data, four unique clusters were identified, each representing different combinations of financial stability and reporting consistency. To capture temporal variations and uncover key behavioral insights, Dynamic Time Warping (DTW), K-Means, and Hierarchical Clustering methods were employed. The results highlight that insurers with stable financial performance and consistent reporting practices emerge as reliable options for policyholders. Conversely, clusters containing underperforming companies and those with reporting gaps reveal significant operational challenges and data consistency issues. These findings underscore the critical role of transparency and timely financial reporting in maintaining the sector’s resilience. This study contributes to the literature by integrating time series analysis into financial clustering, offering practical recommendations for enhancing data governance and financial stability in the insurance industry. Future research could further explore non-financial indicators and alternative clustering approaches to gain deeper insights into performance dynamics.
@@ 1   Introduction
Traditional methods for assessing financial health in the insurance industry often rely on individual financial ratios, which, while useful, fail to capture broader trends and interdependencies among key financial indicators. These static approaches may overlook crucial temporal dynamics that influence an insurer’s long-term viability. Clustering techniques offer a data-driven approach to grouping insurance companies based on their financial performance, revealing hidden patterns and emerging risks that might otherwise go unnoticed. By incorporating time-series analysis, this study enhances the ability to detect financial vulnerabilities, identify high-performing firms, and provide actionable insights for regulators, insurers, and policyholders.
This study applies clustering techniques, particularly Dynamic Time Warping (DTW) and Long Short-Term Memory (LSTM) networks, to analyze insurance firms based on financial ratios derived from publicly available data. The use of temporal clustering of quarterly time-series data allows for a more comprehensive evaluation of financial health, enabling comparisons over time and distinguishing firms based on their financial stability and reporting practices. The findings of this study have implications for multiple stakeholders. Policymakers can leverage the insights to improve sectoral efficiency through regulatory reforms, insurers can identify strategic opportunities for improvement, and policyholders can make informed decisions when selecting providers. The study also contributes to the broader discourse on the intersection of financial performance, risk management, and sustainability in the insurance sector.
This paper is structured as follows: Section 2 provides an overview of the relevant literature. Section 3 details the methodology, including the research design, data sources, financial ratio selection, application of temporal feature fusion, clustering analysis, and cluster evaluation metrics. Section 4 reports the findings, followed by Section 5, which discusses the key results, addresses the study’s limitations, and proposes avenues for future research. Finally, Section 6 concludes the paper. All analyses were performed in Python, using a variety of statistical and machine-learning libraries.
@@ 2   Literature Review
Health insurance plays a critical role in healthcare systems, particularly in developing countries like Kenya, where financial protection against rising medical costs is essential. Despite its significance, health insurance penetration in Kenya remains low, with the Insurance Regulatory Authority (IRA) reporting a penetration rate of only 2.25% in 2021. Insurance density stood at 4,777 KES per capita, indicating limited participation in private health insurance programs. The Kenyan insurance sector, predominantly led by private providers, struggles with extending coverage to low-income and rural populations. These challenges are exacerbated by a high reliance on out-of-pocket payments, further straining financial security at the household level. While much of the existing research examines the financial health and operational efficiency of insurance providers, there remains a lack of comprehensive studies addressing the financial performance of health insurers in Kenya and their contribution to overall market stability.
Prior studies on insurance markets have predominantly focused on traditional financial performance measures such as return on assets (ROA), return on equity (ROE), solvency ratios, and claims ratios. These metrics provide valuable insights into an insurer’s profitability, stability, and risk exposure. For instance, studies such as [6, 18, 20, 21, 22] analyze profitability trends in Kenya’s insurance sector using these indicators. However, these studies primarily assess financial ratios in isolation, rather than exploring their interrelationships or how they collectively shape financial sustainability. The fragmented nature of such analyses limits the ability to capture the dynamic behavior of insurers over time. In contrast, clustering techniques offer a more holistic approach by identifying financial patterns across multiple insurers, enabling segmentation based on shared financial characteristics. Despite the growing use of clustering in financial analysis, its application in assessing insurers’ financial health, particularly within Kenya’s health insurance sector, remains largely unexplored.
Several studies outside the Kenyan context have applied clustering techniques to analyze financial performance. [44] employed fuzzy clustering to group companies based on financial ratios, highlighting common financial characteristics among firms with similar performance. Similarly, [10] used cluster analysis to classify firms by financial strategies, revealing distinct differences between high-tech and basic industries. [16] extended this approach to food retail companies in Hungary and Romania, demonstrating the influence of methodological choices, such as K-Means versus K-Medoids on classification outcomes. [37] introduced a financial K-Means algorithm for classifying companies in Zhejiang province, showing that clustering can be instrumental in assessing loan eligibility. While these studies underscore the utility of clustering in financial performance evaluation, they are largely concentrated in developed economies and industries outside the insurance sector. There remains an evident gap in adapting such methodologies to health insurance markets in sub-Saharan Africa, where financial performance is influenced by unique market dynamics, regulatory environments, and economic constraints.
A notable limitation of prior clustering applications in finance is their reliance on static, one-time financial data rather than time-series financial ratios. The absence of a longitudinal approach makes it difficult to assess how insurers’ financial health evolves over time. Dynamic time-warping and advanced clustering techniques could provide deeper insights into temporal financial patterns, yet studies employing these methodologies in the insurance sector are scarce. Moreover, while existing research emphasizes profitability and solvency, limited attention has been given to how these financial ratios collectively influence the overall health of insurance markets in developing economies. The few studies conducted in Kenya, such as [20, 21, 22, 30], focus on select aspects of financial performance, such as claims incurred ratios and liquidity risk, but do not integrate a comprehensive analysis linking multiple financial indicators to market sustainability.
The theoretical foundation for financial ratio analysis in insurance markets has been predominantly derived from financial distress prediction models and market efficiency theories. Traditional financial distress models, such as Altman’s Z-score [4] and Merton’s Model [28], provide a basis for evaluating insurer solvency and predicting financial instability. Market efficiency theories suggest that financial ratios reflect underlying firm performance and risk exposure, influencing investor confidence and market stability. However, these models have rarely been adapted to explore financial clustering in insurance markets. By integrating clustering methodologies with financial distress prediction models, this study seeks to provide a novel framework for evaluating insurers’ financial health. The application of advanced clustering techniques can enhance understanding of financial stability trends, identifying insurers at risk of distress and offering insights into market competitiveness.
Despite contributions from past research, there remain significant gaps in the literature. First, existing studies on financial performance in the Kenyan insurance market have not examined how financial ratios interrelate to influence overall market health. Most research has analyzed individual aspects, such as profitability or solvency, without considering how these dimensions interact within health insurance firms. Second, there is a lack of studies employing time-series clustering techniques to analyze financial performance evolution among insurers. Current approaches rely on static ratio assessments, which fail to capture long-term financial trajectories. Third, while international studies have demonstrated the potential of clustering for financial classification, its applicability to Kenya’s health insurance sector remains unexplored. The limited penetration of health insurance, coupled with regulatory and economic challenges unique to Kenya, necessitates a localized approach to financial performance analysis. Finally, critical financial indicators such as the Claims Payout Ratio, which reflects an insurer’s efficiency in handling claims, have not been adequately examined in the Kenyan context. Given the pivotal role of claims management in the sustainability of health insurers, understanding its interplay with other financial ratios is essential for evaluating market stability. This study addresses these gaps by employing a clustering approach to classify health insurers based on financial performance patterns, offering a more nuanced perspective on market health than traditional financial ratio analysis alone. By leveraging time-series data and advanced clustering methodologies, this research extends prior applications of financial clustering and contributes to the broader discourse on financial performance assessment in emerging insurance markets.
@@ 3.1   Research Design
The research design integrates statistical and machine learning methods to extract meaningful insights from the data. Techniques such as Long Short-Term Memory (LSTM) for temporal feature fusion, Dynamic Time Warping (DTW) for analyzing temporal trends, and K-Means clustering for identifying patterns were employed to examine the financial performance of insurance providers. These methods were implemented using Python’s object-oriented programming framework, taking advantage of Scikit Learn and clustering algorithms. In addition, several open-source libraries were used to visualize financial ratios, trends, and other relevant metrics. The complete code for this study is publicly accessible on GitHub [49], promoting further development of the analysis framework.
@@ 3.2   Data
These data, covering multiple fiscal years, were sourced directly from the IRA website [1], ensuring both accuracy and adherence to regulatory standards. No additional data collection methods, such as surveys or interviews, were utilized. To maintain data integrity, any missing values for companies were systematically recorded as zero. Comprehensive data preprocessing was conducted to ensure consistency and quality, making the dataset well-suited for robust statistical analysis.
@@ 3.3   Financial Ratios
This study employs key financial ratios to assess the performance and financial health of health insurance providers in Kenya. The selection of these ratios is based on established financial and actuarial research, aligning with prior studies on insurance firm performance, profitability, and risk exposure as reported in [1]. The chosen ratios provide a comprehensive assessment of operational efficiency, market position, and underwriting performance, which are critical in evaluating the sustainability and competitiveness of insurance providers. A critical measure in this analysis is the Expense Ratio, which reflects the proportion of an insurer’s operational costs relative to its income. It is calculated as the ratio of expenses to gross premium income. The expenses are estimated using the formula:
@@
In addition to the Expense Ratio, the study evaluates other financial ratios, including Market Share, Claims Paid Ratio, Claims Incurred Ratio, Claims Payout Ratio, Underwriting Profit Ratio, and the Combined Ratio, as detailed in Table 2. Extreme values were retained in the analysis to ensure meaningful variations, particularly those indicative of financial distress, are not suppressed. The financial ratios were normalized using the StandardScaler, standardizing each feature to have a mean of zero and a standard deviation of one. A within-company scaling approach was used, where scaling parameters were computed separately for each company to preserve unique statistical characteristics.
Ratio Formula Global Range
Market Share PremiumsTotal Market $\frac{\text{Premiums}}{\text{Total Market Premiums}}\times 100$ 5% - 30%
Claims Paid Ratio Claims PaidGross Premium $\frac{\text{Claims Paid}}{\text{Gross Premium Income}}\times 100$ 50% - 80%
Loss Ratio Claims IncurredGross Premium $\frac{\text{Claims Incurred}}{\text{Gross Premium Income}}\times 100$ 60% - 75%
Underwriting Profit Ratio Underwriting ProfitGross Premium $\frac{\text{Underwriting Profit}}{\text{Gross Premium Income}}\times 100$ 5% - 15%
Expense Ratio ExpensesGross Premium $\frac{\text{Expenses}}{\text{Gross Premium Income}}\times 100$ 10% - 25%
@@
Claims Payout Ratio Claims PaidClaims $\frac{\text{Claims Paid}}{\text{Claims Incurred}}\times 100$ 60% - 90%
Table 2: Key Financial Ratios used to evaluate insurance providers. The global range represents typical industry standards for each ratio, indicating the desired financial performance for insurance companies.
@@ 3.4   Long Short-Term Memory (LSTM)
Traditional financial analysis often relies on handcrafted ratios to capture key aspects of a firm’s performance. However, financial data is inherently sequential, with past values influencing future performance. The Long Short-Term Memory (LSTM) model is well-suited for capturing such dependencies by maintaining a ”memory” of relevant information while filtering out noise [13, 43, 26, 39]. Unlike traditional statistical techniques that assume linearity, LSTM learns complex temporal patterns that may not be explicitly defined in financial formulas. In this study, we leverage LSTM not for prediction, but for dimensionality reduction, compressing financial signals into a single latent representation that preserves key financial dynamics over time.
The LSTM architecture is a specialized recurrent neural network (RNN) designed to capture long-range dependencies by addressing the vanishing gradient problem through gating mechanisms. Each LSTM memory block consists of a cell state $\mathbf{c}_{i,j}$ that stores long-term information and three gates, input, forget, and output, that regulate information flow. Given an input sequence $\mathbf{R}\in\mathbb{R}^{N\times J\times F}$ consisting of financial records from $N$ insurance companies over $J$ time points with $F$ financial ratios per time step as introduced in Table 2. The goal is to reduce $F$ to a single temporal dimension $F^{\prime}$, resulting in a compressed representation $\mathbf{Z}\in\mathbb{R}^{N\times J\times F^{\prime}}$. The LSTM memory block operates as follows:
  • Block Input: The block input $\mathbf{g}_{i,j}\in\mathbb{R}^{y}$ combines the current input $\mathbf{r}_{i,j}$ and the previous hidden state $\mathbf{y}_{i,j-1}$:
    where $\mathbf{W}_{g}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{g}\in\mathbb{R}^{y\times y}$, and $\mathbf{b}_{g}\in\mathbb{R}^{y}$ are learnable parameters.
  • Input Gate: Determines the extent to which $\mathbf{g}_{i,j}$ influences the cell state:
    where $\mathbf{W}_{i}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{i}\in\mathbb{R}^{y\times y}$, $\mathbf{b}_{i}\in\mathbb{R}^{y}$, and $\mathbf{p}_{i}\in\mathbb{R}^{y}$ are learnable parameters.
  • Forget Gate: Determines which information from the previous cell state is retained:
    where $\mathbf{W}_{f}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{f}\in\mathbb{R}^{y\times y}$, $\mathbf{b}_{f}\in\mathbb{R}^{y}$, and $\mathbf{p}_{f}\in\mathbb{R}^{y}$ are learnable parameters.
  • Cell State Update: Updates the cell state based on retained and new information:
    $\mathbf{c}_{i,j}=\mathbf{f}_{i,j}\odot\mathbf{c}_{i,j-1}+\mathbf{i}_{i,j}\odot\mathbf{g}_{i,j}.$
  • Output Gate: Controls the information passed to the final output:
    where $\mathbf{W}_{o}\in\mathbb{R}^{y\times F}$, $\mathbf{R}_{o}\in\mathbb{R}^{y\times y}$, $\mathbf{b}_{o}\in\mathbb{R}^{y}$, and $\mathbf{p}_{o}\in\mathbb{R}^{y}$ are learnable parameters.
  • Hidden State Update: Computes the updated hidden state:
    $\mathbf{y}_{i,j}=\tanh(\mathbf{c}_{i,j})\odot\mathbf{o}_{i,j}.$
  • LSTM Output: A fully connected dense layer maps the hidden state to the output:
    $\mathbf{z}_{i,j}=\mathbf{W}_{z}\mathbf{y}_{i,j}+\mathbf{b}_{z},$
    where $\mathbf{W}_{z}\in\mathbb{R}^{1\times y}$ and $\mathbf{b}_{z}\in\mathbb{R}^{1}$ are learnable parameters.
The overall transformation performed by the LSTM network can be expressed as:
$\mathbf{Z}=\psi_{\text{LSTM}}(\mathbf{R}),$
where $\psi_{\text{LSTM}}$ denotes the LSTM’s mapping from the input space $\mathbb{R}^{N\times J\times F}$ to the output space $\mathbb{R}^{N\times J\times F^{\prime}}$. Figure 2 displays the LSTM architecture.
@@
Unlike static dimensionality reduction methods such as PCA or t-SNE, which do not account for temporal dependencies, LSTM dynamically transforms financial data while preserving its sequential nature. By applying LSTM, we extract a latent representation that summarizes evolving financial trajectories while retaining meaningful variations. This compressed representation ensures that firms with similar financial evolution are mapped closer together in latent space, preserving patterns that static methods may overlook.
@@
where $\hat{\mathbf{R}}$ is the reconstructed input tensor. The dataset used in this study consists of 28 insurance companies, each observed over 41 time points with 7 financial ratios (as detailed in Table 2). The LSTM model was applied to extract a lower-dimensional representation, reducing the input tensor $\mathbf{R}\in\mathbb{R}^{28\times 41\times 7}$ to a more compact form, $\mathbf{Z}\in\mathbb{R}^{28\times 41\times 1}$. This transformation preserves essential financial patterns while maintaining temporal dependencies. The extracted latent features, visualized in Fig. 3, serve as the foundation for the clustering analysis discussed in Section 3.5.
Figure 3: The heatmap displays the reduced time series for each company after applying temporal feature fusion using LSTM, which transforms the data from $\mathbb{R}^{7}$ to $\mathbb{R}^{1}$. Color intensity indicates the magnitude of the reduced data values, with darker shades representing higher values. The x-axis represents time points, and the y-axis lists the companies, which will be clustered based on this dataset.